OpenStudy (donnie1999):

HELP WILL MEDAL!!! simplify the expression 64^1/2

Nnesha (nnesha):

you can convert exponent to root example \[\huge\rm x^\frac{ m }{ n } =\sqrt[n]{x^m}\] let m=1 and n=2 n=root number ask if u hve question